if the cube roots of unity are 1,w,w^2then fnd the roots of the equation (x-1)^3+8=0

Grade:

1 Answers

Aditya Gupta
3 years ago
write (x – 1)^3= – 8= – (2)^3
or (1 – x)^3= 2^3
or [(1 – x)/2]^3= 1
or (1 – x)/2= 1^(1/3)= 1, w, w^2
or 1 – x= 2, 2w, 2w^2
or x= – 1, 1 – 2w, 1 – 2w^2.
KINDLY APPROVE :))